A 2.00-kg block is given an initial speed of 2.50 m/s up a plane inclined 15.0° with the horizontal.

The coefficient of kinetic friction between block and plane is 0.250. What speed does it have just as

it returns to its initial position?


1
Expert's answer
2022-01-25T16:05:22-0500

"\frac {mv^2}2=mgh+F_{fr}l,"

"\frac{v^2}2=gh+\frac{\mu gh}{\tan\alpha},"

"h=\frac{v^2\tan\alpha}{2g(\tan\alpha+\mu)};"


"mgh=\frac{mv_1^2}2+Fl,"

"gh=\frac{v_1^2}2+\frac{\mu gh}{\tan\alpha},"

"v_1=v\sqrt{\frac{\tan\alpha-\mu}{\tan\alpha+\mu}}=0.47~\frac ms."